# Hilbert space fragmentation from lattice geometry

### v1

@@ III.3 The modified two-dimensional lattice
The modified two-dimensional lattice is structurally similar to the Vicsek lattice as it is constructed from four first generation Vicsek fractals. It has both sites with four nearest neighbors, which may act as blockades, and sites with two nearest neighbors that do not restrict the movement of domain walls. However, in contrast to the Vicsek fractal, the modified two-dimensional lattice introduces a “loop”. Figure 3(d) shows the Hamiltonian operator for each symmetry sector on the modified two-dimensional lattice. The loop reduces the amount of fragmentation in the modified two-dimensional lattice compared to the Vicsek fractal lattice. For $n_{\text{dw}}\geq 16$, no fragmentation is present. The presence of loops in the two-dimensional lattice and hexaflake lattice might similarly explain the lack of fragmentation on these lattices in symmetry sectors with a large number of domain walls.
@@ IV The autocorrelation function of local observables
In the one-dimensional, scarred model, the time-averaged autocorrelation function converges to a value close to zero. This behavior is expected for a model where the majority of energy eigenstates are thermal. On the four lattices with dimensions larger than one, the Mazur bound is larger than zero, and, consequently, the autocorrelation function converges to a finite value. Hence, the model displays nonthermal behavior on these lattices and the system retains some memory of its initial state. We remark that the Mazur bound is not tight for the considered lattices with dimensions larger than one. However, in appendix D, we show that the Mazur bound becomes tight when the Hamiltonian is perturbed in each Krylov subspace by a random matrix drawn from the Gaussian orthogonal ensemble.
@@ V Conclusion
We generalized a known, one-dimensional, scarred model to higher dimensional lattices. The model displays HSF on the Vicsek fractal lattice of arbitrary dimension, with the Hilbert space shattering into an exponential number of one-dimensional Krylov subspaces. We demonstrated that the largest Krylov subspace constitutes a vanishingly small fraction of the symmetry sector when the number of domain walls is close to zero or close to being maximal. Therefore, the model is strongly fragmented on the Vicsek fractal lattice in these symmetry sectors. Our methods did not allow the analysis of symmetry sectors with the number of domain walls close to half-filling. The model displays features similar to HSF on a section of the second-generation hexaflake fractal lattice, the two-dimensional lattice, and a modified two-dimensional lattice. We studied the time-averaged autocorrelation function of the $z$-component of a single spin and compared it with the corresponding Mazur bound. We demonstrated that the model displays nonthermal dynamics on all the considered lattices with dimensions larger than one by observing a finite value of the long-time average of the autocorrelation function. We also showed that the Mazur bound becomes tight when the Hamiltonian is perturbed by a block diagonal random matrix. This work demonstrates that Hilbert space fragmentation may arise solely from the conservation of the number of domain walls in dimensions larger than one and that the nature of the fragmentation depends on the lattice geometry.
@@ Appendix B Construction of an exponential number of Krylov subspaces on the Vicsek fractal lattice
We prove that the model in Eq. (2) displays Hilbert space fragmentation on the Vicsek fractal lattice of arbitrary generation $g$ by explicitly constructing a set of Krylov subspaces. We focus on one-dimensional Krylov subspaces and demonstrate that the number of such subspaces scales exponentially with the system size.
@@
Consider the Vicsek fractal lattice of generation one. There exist several “frozen states” where all spins are dynamically inactive. We aim to construct one-dimensional Krylov subspaces for generation $g>1$ by utilizing the frozen states from generation one. In the following, we consider the Vicsek fractal lattice of generation $g$ to consist of $5^{g-1}$ generation one fractal lattices. We also consider two generation one fractals to be nearest neighbors if they contain spins that are nearest neighbors. A generation one fractal has one, two, or four other generation one fractals as nearest neighbors as illustrated in Fig. 6. In particular, there are $5^{g-2}$ generation one fractals with four nearest neighbors. On these generation one fractals, we choose all sites to be spin-down. The remaining $4\cdot 5^{g-2}$ generation one fractals have either one or two neighbors, and they allow at least three frozen configurations as illustrated in Fig. 6. We construct one-dimensional Krylov subspaces as the tensor product of frozen states on these generation one fractals. There are at least $3^{4\cdot 5^{g-2}}=(3^{4/25})^{N}$ such one-dimensional Krylov subspaces where $N=5^{g}$ is the total number of sites. The number of Krylov subspaces, hence, grows exponentially with system size and the model from Eq. (2) on the Vicsek fractal lattice displays Hilbert space fragmentation.
@@ Appendix C Estimation of sector sizes and Krylov subspace dimensions
Appendix D Perturbing the Hamiltonian with a random matrix

### v2

@@ III.3 The hexaflake fractal lattice and the modified two-dimensional lattice
The modified two-dimensional lattice is structurally similar to the Vicsek fractal lattice as it is constructed from four first-generation Vicsek fractal lattices. It consists of sites with four nearest neighbors which may act as blockades and sites with two nearest neighbors that do not restrict the movement of the domain walls. However, in contrast to the Vicsek fractal lattice, the modified two-dimensional lattice introduces a “loop”. Figure 3(d) illustrates the Krylov subspaces for each symmetry sector on the modified two-dimensional lattice. We observe that the model displays a smaller amount of fragmentation on this lattice as compared to the second-generation Vicsek fractal lattice. In particular, symmetry sectors with $n_{\text{dw}}\geq 16$ domain walls display no fragmentation. The reduced amount of fragmentation may be related to the presence of a loop in the modified two-dimension lattice. Similarly, the presence of loops in the two-dimensional lattice and hexaflake fractal lattice may be related to the reduced fragmentation of the model on these lattices in symmetry sectors with a large number of domain walls.
We emphasize that the results presented in this section concern a single system size for the hexaflake fractal lattice and the modified two-dimensional lattice and the observed behavior may not generalize to larger system sizes.
@@ IV The autocorrelation function of local observables
In the one-dimensional, scarred model, the time-averaged autocorrelation function converges to a value close to zero. This behavior is expected for a model where the majority of energy eigenstates are thermal. On the four lattices with dimensions larger than one, the Mazur bound is larger than zero, and, consequently, the autocorrelation function converges to a finite value. Hence, the model displays nonthermal behavior on these lattices and the system retains some memory of its initial state. We remark that the Mazur bound is not tight for the considered lattices with dimensions larger than one. However, in appendix E, we show that the Mazur bound becomes tight when the Hamiltonian is perturbed in each Krylov subspace by a random matrix drawn from the Gaussian orthogonal ensemble.
@@ V Conclusion
We generalized a known, one-dimensional, scarred model to higher dimensional lattices. The model displays HSF on the Vicsek fractal lattice and the two-dimensional lattice, with the Hilbert space shattering into an exponential number of one-dimensional Krylov subspaces for both lattices. For the Vicsek fractal lattice, we demonstrated that the largest Krylov subspace constitutes a vanishingly small fraction of the symmetry sector when the number of domain walls is close to zero or close to being maximal. Therefore, the model is strongly fragmented on the Vicsek fractal lattice in these symmetry sectors. For the two-dimensional lattice, the model is strongly fragmented for a low density of domain walls and weakly fragmented for a high density of domain walls. For the Vicsek fractal lattice and the two-dimensional lattice, we also showed that the fragmentation persists at a finite domain wall density in the thermodynamic limit. The model displays features similar to HSF on a section of the second-generation hexaflake fractal lattice and a modified two-dimensional lattice. We studied the time-averaged autocorrelation function of the $z$-component of a single spin and compared it with the corresponding Mazur bound. We demonstrated that the model displays nonthermal dynamics on all the considered lattices with dimensions larger than one by observing a finite value of the long-time average of the autocorrelation function. We also showed that the Mazur bound becomes tight when the Hamiltonian is perturbed by a block diagonal random matrix. This work demonstrates that Hilbert space fragmentation may arise solely from the conservation of the number of domain walls in dimensions larger than one and that the nature of the fragmentation depends on the lattice geometry.
@@ Appendix B Construction of an exponential number of Krylov subspaces on the Vicsek fractal lattice
We prove that the model in Eq. (2) displays Hilbert space fragmentation on the Vicsek fractal lattice by explicitly constructing a set of Krylov subspaces. We focus on one-dimensional Krylov subspaces and demonstrate that the number of such subspaces scales exponentially with the system size. We also show that the fragmentation persists at a finite domain wall density in the thermodynamic limit.
@@
Consider the Vicsek fractal lattice of generation one. There exist several “frozen states” where all spins are dynamically inactive. We aim to construct one-dimensional Krylov subspaces for generation $g>1$ by utilizing the frozen states from generation one. In the following, we consider the Vicsek fractal lattice of generation $g$ to consist of $5^{g-1}$ first-generation fractal lattices. We also consider two first-generation fractals to be nearest neighbors if they contain spins that are nearest neighbors. A first-generation fractal has one, two, or four other first-generation fractals as nearest neighbors as illustrated in Fig. 7. In particular, there are $5^{g-2}$ first-generation fractals with four nearest neighbors. On these first-generation fractals, we choose all sites to be spin-down. The remaining $4\cdot 5^{g-2}$ first-generation fractals have either one or two nearest neighbors, and they allow at least three frozen configurations as illustrated in Fig. 7. We construct one-dimensional Krylov subspaces as the tensor product of frozen states on these first-generation fractals. There are at least $3^{4\cdot 5^{g-2}}=3^{4N/25}$ such one-dimensional Krylov subspaces where $N=5^{g}$ is the total number of sites. The number of Krylov subspaces, hence, grows exponentially with system size, and the model from Eq. (2) displays Hilbert space fragmentation on the Vicsek fractal lattice.
Next, we show that the fragmentation persists at a finite density of domain walls in the thermodynamic limit. Similarly to the previous paragraph, we consider the Vicsek fractal lattice of generation $g$ to consist of $5^{g-1}$ first-generation Vicsek fractal lattices. The $4\cdot 5^{g-2}$ first-generation fractals with one or two nearest neighbors may be chosen from at least three frozen spin configurations as illustrated in Fig. 7. One of these configurations does not introduce any domain walls since all sites are spin-down. The remaining configurations have one spin-up which introduces two domain walls. By choosing $m$ first-generation fractals as a frozen configuration with a single spin-up and all other first-generation fractals as spin-down, we construct a frozen state with $n_{\text{dw}}=2m$ domain walls where $m=0,1,\ldots,4\cdot 5^{g-2}$. The number of ways to choose the $m$ first-generation fractal lattices is given by the binomial coefficient (${4\cdot 5^{g-2}\choose m}$. Furthermore, for each first-generation fractal, there are at least two frozen configurations with a single spin-up. Therefore, the number of frozen states with $n_{\text{dw}}=2m$ domain walls is at least
(${4N/25\choose n_{\text{dw}}/2}2^{n_{\text{dw}}/2}.$ (11)
where $N=5^{g}$ is the number of sites. We aim to determine the number of frozen states at a fixed density of domain walls $\rho_{\text{dw}}=n_{\text{dw}}/n_{\text{dw}}^{\text{max}}$ where $n_{\text{dw}}^{\text{max}}$ is the maximum number of domain walls on the Vicsek fractal lattice. Before rewriting Eq. (11) in terms of the domain wall density $\rho_{\text{dw}}$, we determine the maximum number of domain walls $n_{\text{dw}}^{\text{max}}$.
The Vicsek fractal lattice is bipartite and the maximum number of domain walls is obtained by choosing all sites in one part as spin-down and all sites in the other part as spin-up. In this case, all nearest neighbor edges are occupied by a domain wall. Hence, the maximum number of domain walls is equal to the number of nearest neighbor edges. In the following, we explicitly include the generation $g$ in the notation, i.e., $n_{\text{dw},g}^{\text{max}}$ is the maximum number of domain walls on the Vicsek fractal lattice of generation $g$. We determine the number of nearest neighbor edges recursively. The Vicsek fractal lattice of generation $g+1$ is constructed by connecting five Vicsek fractal lattices of generation $g$ as illustrated in Figs. 1(a)-(c). The number of nearest neighbor edges for generation $g+1$ is, therefore, five times that for generation $g$ minus four corresponding to where the fractals of generation $g$ are connected. The maximum number of domain walls follows the same relationship
$n_{\text{dw},g+1}^{\text{max}}=5n_{\text{dw},g}^{\text{max}}-4.$ (12)
The solution to the recursive relation with the initial condition $n_{\text{dw},g=1}^{\text{max}}=8$ is given by
$n_{\text{dw},g}^{\text{max}}=7\cdot 5^{g-1}+1$ (13)
Inserting Eq. (13) and $\rho_{\text{dw}}=n_{\text{dw}}/n_{\text{dw}}^{\text{max}}$ into Eq. (11), we obtain a lower bound on the number of frozen states at domain wall density $\rho_{\text{dw}}$
(${4N/25\choose\rho_{\text{dw}}(7N/5+1)/2}2^{\rho_{\text{dw}}(7N/5+1)/2}$ (14)
Recall that the number of domain walls for the considered frozen states is given by $n_{\text{dw}}=2m$ with $m=0,1,\ldots,4\cdot 5^{g-2}$. Therefore, the expression in (14) is valid for all domain wall densities up to
$\rho_{\text{dw}}$ $=\frac{8\cdot 5^{g-2}}{7\cdot 5^{g-1}+1}\xrightarrow[g\to\infty]{}\frac{8}{35}.$ (15)
Equations (14) and (15) demonstrate that the fragmentation persists at all domain wall densities $\rho_{\text{dw}}\leq 8/35$ in the thermodynamic limit since the number of Krylov subspaces scales exponentially with system size at these domain wall densities.
@@ Appendix C Estimation of sector sizes and Krylov subspace dimensions
Appendix D Construction of an exponential number of Krylov subspaces on the two-dimensional lattice
Figure 8: The two-dimensional lattice $\mathcal{L}$ of size $L_{x}\times L_{y}$ (light red and blue circles). The lattice is padded with extra sites along its boundary (dark red circles) to ensure all sites in the original lattice have an even number of nearest neighbors. The extra sites are dynamically inactive and we choose them to be spin-down. The lattice $\mathcal{L}$ does not contain the dynamically inactive sites. The sublattice $\mathcal{L}^{\prime}\subset\mathcal{L}$ (blue circles) form a $5$-periodic pattern on $\mathcal{L}$ in the $x$ and $y$ directions. The lattice $\mathcal{L}$ is separated into four parts $\mathcal{L}_{A}$, $\mathcal{L}_{B}$, $\mathcal{L}_{C}$, and $\mathcal{L}_{D}$ (gray boxes). Part $\mathcal{L}_{A}$ is a rectangular lattice where both sides are a multiple of $5$. Part $\mathcal{L}_{B}$ contains no sites when $\operatorname{mod}_{5}(L_{x})=0$ and $\mathcal{L}_{C}$ contains no sites when $\operatorname{mod}_{5}(L_{y})=0$. When parts $\mathcal{L}_{B}$ and $\mathcal{L}_{C}$ are nonempty, they are rectangular lattices where one side is a multiple of $5$. Part $\mathcal{L}_{D}$ is empty when $\operatorname{mod}_{5}(L_{x})=0$ or $\operatorname{mod}_{5}(L_{y})=0$. When part $\mathcal{L}_{D}$ is not empty it is a rectangular lattice where neither side is a multiple of $5$.
We demonstrate that the model in Eq. (2) displays HSF on the two-dimensional lattice by constructing an exponential number of one-dimensional Krylov subspaces. We consider the two-dimensional lattice of size $L_{x}\times L_{y}$ given by
[…]
We remark that the lattice $\mathcal{L}$ only contains the dynamically active sites and, hence, does not contain the extra sites introduced in Sec. II to ensure all sites in the original lattice have an even number of nearest neighbors. We also consider the sublattice $\mathcal{L}^{\prime}\subset\mathcal{L}$ given by
[…]
Figure 8 illustates the lattice $\mathcal{L}$ and the sublattice $\mathcal{L}^{\prime}$. The sites contained in $\mathcal{L}^{\prime}$ form a pattern on the lattice $\mathcal{L}$ that repeats in the $x$ and $y$ directions with a period of $5$, i.e., if $(x,y)\in\mathcal{L}^{\prime}$ and $0\leq x\pm 5<L_{x}$ then $(x\pm 5,y)\in\mathcal{L}^{\prime}$ and, similarly, if $(x,y)\in\mathcal{L}^{\prime}$ and $0\leq y\pm 5<L_{y}$ then $(x,y\pm 5)\in\mathcal{L}^{\prime}$.
We aim to construct an exponential number of product states that are exact energy eigenstates of the Hamiltonian operator in Eq. (2). We consider product states of the form $\lvert s_{1}s_{2}\ldots s_{L_{x}L_{y}}\rangle$ where $s_{i}\in\{-1/2,1/2\}$ is the spin along the $z$-direction on site $i$. These product states are eigenstates of Eqs. (3b) and (3c) for all values of the spins $\{s_{i}\}_{i=1}^{L_{x}L_{y}}$. We ensure the product states are eigenstates of the full Hamiltonian operator by choosing the spins such that the kinetic term in Eq. (3a) annihilates the product states. We take all sites in $\mathcal{L}\setminus\mathcal{L}^{\prime}$ to be spin-down where $\mathcal{L}\setminus\mathcal{L}^{\prime}$ is the set difference, i.e., the sites contained in $\mathcal{L}$ but not in $\mathcal{L}^{\prime}$. We choose the sites in $\mathcal{L}^{\prime}$ freely as either spin-down or spin-up. Recall that the kinetic term in Eq. (3a) flips a spin if the sum of the magnetization of its nearest neighbors is zero. When choosing the spins as described above, the sum of the magnetization of the nearest neighbors is either $-2$ or $-1$ for all sites. Therefore, all product states of this form are eigenstates of the full Hamiltonian operator. Since the sites in $\mathcal{L}^{\prime}$ may be chosen as either spin-down or spin-up, the number of such product states is given by $2^{\lvert\mathcal{L}^{\prime}\rvert}$ where $\lvert\mathcal{L}^{\prime}\rvert$ is the number of sites in $\mathcal{L}^{\prime}$.
[…]
$\mathcal{L}_{A}$ =$=\big{\{}(x,y)\in\mathcal{L}\mid x<5\lfloor L_{x}/5\rfloor,\,y<5\lfloor L_{y}/5\rfloor\big{\}},$ (23)
$\mathcal{L}_{B}$ =$=\big{\{}(x,y)\in\mathcal{L}\mid 5\lfloor L_{x}/5\rfloor\leq x,\,y<5\lfloor L_{y}/5\rfloor\big{\}},$ (24)
[…]
We also define the sites from $\mathcal{L}^{\prime}$ in each of these parts by $\mathcal{L}_{X}^{\prime}=\mathcal{L}^{\prime}\cap\mathcal{L}_{X}$ with $X\in\{A,B,C,D\}$. The number of sites in $\mathcal{L}^{\prime}$ is given by
$\lvert\mathcal{L}^{\prime}\rvert=\lvert\mathcal{L}_{A}^{\prime}\rvert+\lvert\mathcal{L}_{B}^{\prime}\rvert+\lvert\mathcal{L}_{C}^{\prime}\rvert+\lvert\mathcal{L}_{D}^{\prime}\rvert.$ (27)
Part $\mathcal{L}_{A}$ represents a rectangular lattice with sides $5\lfloor L_{x}/5\rfloor$ and $5\lfloor L_{y}/5\rfloor$ where both side lengths are a multiple of $5$. Recall that the sites in $\mathcal{L}^{\prime}$ form a pattern on the lattice $\mathcal{L}$ with period $5$ along the $x$ and $y$ directions. Therefore, any horizontal sequence of five sites in $\mathcal{L}$, i.e., $(x,y),(x+1,y),\ldots(x+4,y)\in\mathcal{L}$, contains exactly one site in $\mathcal{L}^{\prime}$. Similarly, any vertical sequence of five sites in $\mathcal{L}$ also contain exactly one site in $\mathcal{L}^{\prime}$. Hence, the number of sites in $\mathcal{L}_{A}^{\prime}$ is given by
$\lvert\mathcal{L}_{A}^{\prime}\rvert=5\left\lfloor\frac{L_{x}}{5}\right\rfloor\left\lfloor\frac{L_{y}}{5}\right\rfloor.$ (28)
Part $\mathcal{L}_{B}$ is a rectangular lattice with side lengths $\operatorname{mod}_{5}(L_{x})$ and $5\lfloor L_{y}/5\rfloor$. Since the length of one side is a multiple of $5$, the number of sites in $\mathcal{L}^{\prime}_{B}$ is given by
$\lvert\mathcal{L}_{B}^{\prime}\rvert=\operatorname{mod}_{5}(L_{x})\left\lfloor\frac{L_{y}}{5}\right\rfloor.$ (29)
Finally, part $\mathcal{L}_{C}$ forms a rectangle with side lengths $\operatorname{mod}_{5}(L_{y})$ and $5\lfloor L_{x}/5\rfloor$. Again, one side length is a multiple of $5$ and the number of sites in $\mathcal{L}_{C}^{\prime}$ is given by
$\lvert\mathcal{L}_{C}^{\prime}\rvert=\left\lfloor\frac{L_{x}}{5}\right\rfloor\operatorname{mod}_{5}(L_{y}).$ (30)
Inserting Eqs. (28)-(30) into Eq. (27), we find
$\lvert\mathcal{L}^{\prime}\rvert=\frac{L_{x}L_{y}}{5}-\frac{1}{5}\operatorname{mod}_{5}(L_{x})\operatorname{mod}_{5}(L_{y})+\lvert\mathcal{L}_{D}^{\prime}\rvert$ (31)
where we have utilized that $5\lfloor L/5\rfloor=L-\operatorname{mod}_{5}(L)$. We consider all possible values of $\operatorname{mod}_{5}(L_{x}),\,\operatorname{mod}_{5}(L_{y})\in\{0,1,2,3,4\}$ and find that the expression $-\operatorname{mod}_{5}(L_{x})\operatorname{mod}_{5}(L_{y})/5+\lvert\mathcal{L}_{D}^{\prime}\rvert$ is always larger than or equal to $-2/5$. Hence, we find the lower bound on the number of sites in $\mathcal{L}^{\prime}$ given by
$\lvert\mathcal{L}^{\prime}\rvert\geq\frac{L_{x}L_{y}-2}{5}$ (32)
In total, we find that the number of sites in $\mathcal{L}^{\prime}$ scales linearly with the system size. Therefore, the number of one-dimensional Krylov subspaces is larger than $2^{(L_{x}L_{y}-2)/5}$ and, hence, grows exponentially with system size. In summary, the model from Eq. (2) displays HSF on the two-dimensional lattice.
We show that the fragmentation persists at a finite density of domain walls in the thermodynamic limit. We follow the same approach as in Appendix B and utilize the frozen configurations constructed in the previous paragraph. We consider product states where all sites are spin-down except the sites in $\mathcal{L}^{\prime}$ which we choose as either spin-down or spin-up. By choosing $m$ sites in $\mathcal{L}^{\prime}$ as spin-up, we construct a frozen state with $n_{\text{dw}}=4m$ domain walls where $m=0,1,\ldots,\lvert\mathcal{L}^{\prime}\rvert$. The binomial coefficient (${\lvert\mathcal{L}^{\prime}\rvert\choose m}$ counts the number of ways of choosing $m$ sites from $\mathcal{L}^{\prime}$. Using the inequality (32), we obtain a lower bound on the number of frozen states with $n_{\text{dw}}$ domain walls
(${(N-2)/5\choose n_{\text{dw}}/4}$ (33)
where $N=L_{x}L_{y}$ is the number of sites. Before rewriting this expression in terms of the domain wall density $\rho_{\text{dw}}=n_{\text{dw}}/n_{\text{dw}}^{\text{max}}$, we determine the maximum number of domain walls $n_{\text{dw}}^{\text{max}}$ on the two-dimensional lattice.
The maximum number of domain walls is obtained by choosing every other site as spin-up and the remaining sites as spin-down. Concretely, we choose the lattice sites ~=$\widetilde{\mathcal{L}}=\{(x,x+2\ell)\in\mathcal{L}\mid x,\ell\in\mathds{Z}\}$ as spin-up and the remaining sites as spin-down. Recall that the two-dimensional lattice is padded with extra sites as illustrated in Fig. 8 and we refer to these sites as the dynamically inactive sites. For the spin configuration where all sites in $\widetilde{\mathcal{L}}$ are spin-up and the remaining sites are spin-down, the number of domain walls between dynamically active sites is given by $2L_{x}L_{y}-L_{x}-L_{y}$. The number of domain walls between the dynamically active sites and the dynamically inactive sites depends on the size of the lattice. If $L_{x}$ and $L_{y}$ are odd, there are $L_{x}+L_{y}+2$ domain walls between these sites. Otherwise, there are $L_{x}+L_{y}$ domain walls. We cover both scenarios by writing the number of domain walls between the dynamically active sites and the dynamically inactive sites according to $L_{x}+L_{y}+2\operatorname{mod}_{2}(L_{x})\operatorname{mod}_{2}(L_{y})$. We use the notation $\mu=\operatorname{mod}_{2}(L_{x})\operatorname{mod}_{2}(L_{y})$ in the following to keep the expressions concise. The maximum number of domain walls is given by
$n_{\text{dw}}^{\text{max}}=2(N+\mu).$ (34)
We determine a lower bound on the number of frozen states at domain wall density $\rho_{\text{dw}}$ by inserting the expression $\rho_{\text{dw}}=n_{\text{dw}}/n_{\text{dw}}^{\text{max}}$ and Eq. (34) into Eq. (33) and using Stirling’s approximation for the factorial $x!\approx\sqrt{2\pi x}x^{x}\exp(-x)$ which is valid in the limit $x\to\infty$
(${(N-2)/5\choose\rho_{\text{dw}}(N+\mu)/2}$
$N\to\infty]{}$ $\frac{1}{\sqrt{2\pi N}\sqrt[10]{5}}\left(\frac{2}{\rho_{\text{dw}}}\right)^{\frac{1+\mu\rho_{\text{dw}}}{2}}\left(\frac{10}{2-5\rho_{\text{dw}}}\right)^{\frac{1-5\mu\rho_{\text{dw}}}{10}}$
×[$\frac{1}{\sqrt[5]{5}}\left(\frac{2}{\rho_{\text{dw}}}\right)^{\frac{\rho_{\text{dw}}}{2}}\left(\frac{10}{2-5\rho_{\text{dw}}}\right)^{\frac{2-5\rho_{\text{dw}}}{10}}\right]^{N}.$ (35)
For the domain wall densities $\rho_{\text{dw}}\in(0,2/5)$, the expression in the square brackets in the third line of Eq. (35) lies in the interval $(1,\sqrt[5]{2})$. Hence, the number of frozen states at a fixed domain wall density scales exponentially with system size for these domain wall densities. The number of domain walls of the considered frozen states is given by $n_{\text{dw}}=4m$ with $m=0,1,\ldots,\lvert\mathcal{L}^{\prime}\rvert$. Using the inequality (32), we find that the frozen states exist at least up to domain wall density
$\rho_{\text{dw}}=\frac{4\lvert{\mathcal{L}}^{\prime}\rvert}{n_{\text{dw}}^{\text{max}}}\geq\frac{2}{5}\frac{N-2}{N+\mu}\xrightarrow[N\to\infty]{}\frac{2}{5}$ (36)
in the thermodynamic limit. Equations (35)-(36) show that the fragmentation persists in the thermodynamic limit at a finite density of domain walls up to at least $\rho_{\text{dw}}=2/5$.
Appendix E Perturbing the Hamiltonian with a random matrix